# Uncertainty-Aware 4D Gaussian Splatting for Monocular Occluded Human Rendering

### v1

@@ 4.3 Double Rasterization
4.4 Optimization via MAP Estimation
To reconstruct high-fidelity humans from monocular videos, we jointly optimize the canonical Gaussian parameters and the deformation network weights . Following the MAP formulation introduced in Sec. 3, our objective function maximizes the posterior . This decomposes the optimization into two synergistic components: a Likelihood term that models data fidelity under heteroscedastic noise, and a Prior term that enforces physical constraints via Confidence-Aware Regularizations.
@@ 4.4 Optimization via MAP Estimation
(2) Uncertainty-Weighted Temporal Inertia: To prevent jittering in the absence of visual cues, we enforce a smoothness prior on the trajectory of uncertain Gaussians. Considering that we train on monocular video sequences with a frame interval , we formulate this constraint as a second-order difference over the sampled timestamps:
@@
Note that in both and , we apply a stop-gradient operator to the uncertainty weights . This ensures that serves purely as an attention mechanism for the regularization strength, preventing the optimizer from trivially reducing to minimize the regularization penalty.
@@
Total Loss. The final objective function consolidates the MAP formulation, combining the uncertainty-weighted likelihood (data fidelity), the confidence-aware priors (regularization), and auxiliary perceptual constraints:
@@
where ensures perceptual quality using standard metrics [wang2004image, zhang2018unreasonable]. Additionally, imposes a global silhouette constraint, aligning the accumulated opacity with the projected SMPL mask to guide coarse geometry. The hyperparameters balance the contribution of each term.
[…]
| Standard Human Rendering | |||||||||||||
| HumanNeRF [weng2022humannerf] | - | 20.67 | 0.9509 | - | - | - | - | ||||||
| GaussianAvatar [hu2024gaussianavatar] | - | 18.01 | 0.9512 | 60.33 | - | - | - | ||||||
| GauHuman [hu2024gauhuman] | - | 21.55 | 0.9430 | 55.88 | 15.09 | 0.8525 | 107.1 | ||||||
| Occlusion-Aware Approaches | |||||||||||||
| OccNeRF [xiang2023rendering] | Scene Decoupling | 22.40 | 0.9562 | 43.01 | 15.71 | 0.8523 | 82.90 | ||||||
| Wild2Avatar [xiang2025rendering] | - | - | - | 14.09 | 0.8484 | 93.31 | |||||||
| OccGaussian [ye2025occgaussian] | Geometric Heuristics | 23.29 | 0.9482 | 41.93 | - | - | - | ||||||
| SymGaussian [jiang2025symgaussian] | 23.22 | 0.9535 | 39.02 | - | - | - | |||||||
| GTU [lee2024guess] | Generative Priors | 22.89 | 0.9503 | 40.78 | 15.83 | 0.8437 | 83.46 | ||||||
| OccFusion [sun2024occfusion] | 23.96 | 0.9548 | 32.34 | 18.28 | 0.8875 | 82.42 | |||||||
| U-4DGS (Ours) | 24.62 | 0.9606 | 31.72 | 20.11 | 0.9030 | 79.17 | |||||||
5 Experiments
5.1 Experimental Setup
Datasets. We evaluate U-4DGS on two standard benchmarks representing both synthetic and real-world occlusion scenarios. First, we utilize the ZJU-MoCap dataset [peng2021neural], which contains multi-view sequences of 6 subjects. Following the protocol of OccNeRF [xiang2023rendering], we simulate occlusions by masking the central 50% of the human region in the first 80% of the frames. We adopt a monocular setting for training, using only Camera 1 with 100 frames subsampled at an interval of 5, while the remaining 22 cameras are reserved for novel view evaluation. Second, to assess robustness in natural environments, we employ the OcMotion dataset [huang2022occluded], which features varying degrees of real-world human-object interactions. Consistent with prior works [xiang2025rendering, sun2024occfusion], we select 6 representative sequences and train on sparse sub-sequences of 50 frames to challenge the model’s capability under limited observations.
Baselines. We benchmark our method against comprehensive state-of-the-art approaches, categorized into two distinct groups: (1) Standard Human Rendering. To demonstrate the impact of occlusion on conventional pipelines, we evaluate HumanNeRF [weng2022humannerf], GaussianAvatar [hu2024gaussianavatar], and GauHuman [hu2024gauhuman]. These methods are trained directly on the occluded sequences without specific handling mechanisms. (2) Occlusion-Aware Approaches. We compare against representative methods covering three mainstream technical paradigms: a) Scene Decoupling: OccNeRF [xiang2023rendering] and Wild2Avatar [xiang2025rendering], which utilize implicit fields or layer-wise parameterization to separate the human subject from environmental obstacles. b) Geometric Heuristics: OccGaussian [ye2025occgaussian] and SymGaussian [jiang2025symgaussian], which rely on rigid priors such as feature aggregation or left-right symmetry to infer missing geometry. c) Generative Priors: OccFusion [sun2024occfusion] and GTU [lee2024guess], which leverage 2D diffusion models to hallucinate the unobserved regions. For a fair comparison, all methods are evaluated using identical segmentation masks and SMPL pose priors.
Implementation Details: Our framework is implemented in PyTorch [paszke2019pytorch] and trained on a single NVIDIA A100 PCIe GPU. The Probabilistic Deformation Network is instantiated as an 8-layer MLP (hidden dimension 256) with a skip connection at the 4-th layer. To capture high-frequency spatial-temporal dynamics, we apply sinusoidal positional encoding to input coordinates and time with frequency bands and , respectively.
We optimize the model for 10k iterations using the Adam optimizer [kingma2014adam]. Learning rates for canonical Gaussian attributes follow the schedule of GauHuman [hu2024gauhuman], while the deformation network utilizes a rate of . Crucially, to preclude trivial solutions (where the model simply predicts infinite uncertainty to minimize the NLL), we adopt a progressive training strategy. For the initial 2k iterations, we deactivate the uncertainty branch and supervise geometry solely via standard loss. Subsequently, we activate the uncertainty head and switch to the MAP objective (Eq. 7) to jointly optimize for reconstruction fidelity and reliability. The regularization weights are set as: , , and .
5.2 Comparisons with State-of-the-Arts
Quantitative Results. Tab. 4.4 reports the performance on ZJU-MoCap and OcMotion benchmarks. On the synthetic ZJU-MoCap dataset, U-4DGS sets a new SOTA, achieving a PSNR of 24.62 dB. Notably, it surpasses the leading generative baseline, OccFusion (23.96 dB), validating that our uncertainty-guided aggregation yields superior fidelity without relying on potentially unstable diffusion priors. Crucially, on the real-world OcMotion dataset, our framework demonstrates exceptional robustness, outperforming the runner-up by a substantial margin of 1.83 dB. Unlike standard rendering or scene decoupling methods that degrade under occlusions, U-4DGS effectively leverages Confidence-Aware Regularizations to shield the canonical geometry, proving that modeling heteroscedastic noise is decisive for handling unconstrained interactions.
Qualitative Results. Fig. 3 visualizes the rendering quality. Standard methods like GauHuman (GH) [hu2024gauhuman] lack occlusion awareness, erroneously fusing environmental obstacles (e.g., bars or boxes) directly onto the human body. While generative approaches like OccFusion (OF) [sun2024occfusion] can remove occluders, they suffer from inherent stochasticity, often yielding over-smoothed textures or hallucinations inconsistent with the subject’s identity. In contrast, U-4DGS achieves superior fidelity. By functioning as an adaptive gradient modulator, our uncertainty mechanism effectively attenuates gradients from artifacts. By strictly aggregating valid temporal information, our method faithfully restores clean, sharp, and physically consistent avatars, even under severe occlusions.
5.3 Ablation and Analysis
Ablation Studies. We validate the efficacy of each component by incrementally integrating it into the baseline to instantiate our complete MAP framework. Quantitative metrics on ZJU-MoCap are reported in Tab. 2, and visual comparisons are presented in Fig. 4. (1) Exp. A (Baseline): We establish a deterministic baseline OccGauHuman, structurally inheriting the pipeline from GauHuman [hu2024gauhuman]. It employs a standard deformation field optimized via loss, effectively assuming homoscedastic observation noise. As shown in Fig. 4 (Exp. A), limited by this rigid deterministic fitting, the model fails to distinguish between the human body and environmental obstacles. Consequently, it indiscriminately minimizes geometric error, leading to severe artifact adhesion at occlusion boundaries. (2) Exp. B (+ Uncertainty Modeling): Replacing the deterministic field with our Probabilistic Deformation Network and enabling Double Rasterization yields a substantial quality leap. This step introduces the Likelihood term of our MAP objective. By functioning as an adaptive gradient modulator, the learned uncertainty effectively attenuates gradients from unreliable observations, successfully removing occlusion artifacts. However, without regularization priors, the geometry in regions lacking reliable visual cues remains noisy and ill-defined (see Exp. B), as the probabilistic loss solely down-weights error without explicitly guiding completion. (3) Exp. C (+ ): The introduction of Confidence-Guided Spatial Consistency imposes the first Prior term. By leveraging uncertainty to selectively propagate spatial validity from confident neighbors, this constraint effectively regularizes the noisy geometry, restoring a coherent and complete body shape. (4) Exp. D (+ ): Finally, incorporating Uncertainty-Weighted Temporal Inertia completes the U-4DGS framework. This temporal prior ensures trajectory smoothness and further refines surface details, achieving the most faithful reconstruction with high physical plausibility.
@@ 5.3 Ablation and Analysis
| B | + Uncertainty Modeling | 24.15 | 0.9575 | 35.20 |
| C | + Spatial Constraint () | 24.48 | 0.9592 | 32.50 |
| D | + Temporal Constraint () | 24.62 | 0.9606 | 31.72 |
@@
Visualization of Uncertainty Maps. To validate the efficacy of our heteroscedastic noise modeling, we visualize the rendered uncertainty map in Fig. 5. In our MAP framework, represents the estimated scale of observation noise, tasked with autonomously distinguishing between valid signals and occlusion artifacts. As evident in Fig. 5, the network assigns low uncertainty (depicted in dark purple) to visible body parts, allowing the photometric supervision to drive the optimization in these trusted regions. Crucially, for anatomical regions that are occluded in the current frame but belong to the canonical topology, the network autonomously predicts significantly high uncertainty (depicted in bright yellow). This visualization empirically confirms that functions as an adaptive gradient modulator. By automatically increasing the denominator in the NLL objective (Eq. 7), it effectively attenuates gradients in these regions lacking reliable visual cues, preventing the canonical geometry from adhering to environmental occlusions.
Temporal Stability Analysis. To validate the temporal coherence of our framework, we conduct a longitudinal comparison against the generative baseline GTU [lee2024guess] across a long sequence (Fig. 6). As observed in the top row, GTU suffers from severe identity drift due to the inherent stochasticity of its diffusion priors. Since the method hallucinates missing regions frame-by-frame without strict temporal coupling, the synthesized textures fluctuate drastically. This instability becomes particularly pronounced in later frames (e.g., to ), where the clothing consistency collapses, degenerating into severe dark artifacts and incoherent noise on the subject’s back. In contrast, U-4DGS achieves rigorous stability. By anchoring appearance in a unified canonical space and enforcing Uncertainty-Weighted Temporal Inertia, our method effectively constrains the optimization trajectory. This ensures that the reconstructed human maintains physical consistency and sharp details throughout the sequence, eliminating the texture flickering observed in generative approaches.
6 Conclusion
@@
In this paper, we present U-4DGS, a robust framework that reformulates monocular human rendering as an MAP estimation problem to handle severe environmental occlusion. Departing from paradigms that rely on stochastic hallucinations or rigid priors, we explicitly model the heteroscedastic nature of observation noise. By integrating a Probabilistic Deformation Network with a Double Rasterization pipeline, our method establishes an adaptive gradient modulator, enabling the optimization to autonomously distinguish between valid geometric cues and occlusion artifacts. Furthermore, we leverage the learned uncertainty to drive Confidence-Aware Regularizations, preventing geometric drift and ensuring structural integrity in regions lacking reliable visual cues. Extensive evaluations on ZJU-MoCap and OcMotion demonstrate that U-4DGS significantly outperforms state-of-the-art approaches, achieving superior robustness and photorealism.

### v2

@@ 4.3. Joint Rasterization
4.4. Optimization via MAP Estimation
To reconstruct high-fidelity humans from monocular videos, we jointly optimize the Gaussian parameters and the deformation network weights . Following the MAP formulation introduced in Sec. 3, our objective function maximizes the posterior . This decomposes the optimization into two synergistic components: a Likelihood term that models data fidelity under heteroscedastic noise, and a Prior term that enforces physical constraints via Confidence-Aware Regularizations.
@@ 4.4. Optimization via MAP Estimation
(2) Uncertainty-Weighted Temporal Inertia: To prevent jittering in the absence of visual cues, we enforce a smoothness prior on the trajectory of uncertain Gaussians. Considering that we train on video sequences with a frame interval , we formulate this constraint as a second-order difference over the sampled timestamps:
@@
In both and , we apply a stop-gradient operator to the uncertainty weights to prevent degenerate optimization. Because scales the regularization penalties, unconstrained backpropagation would encourage the network to trivially drive simply to minimize these losses, invalidating its role in modeling heteroscedastic observation noise. By decoupling the gradient flow, ensures that the aleatoric uncertainty is learned exclusively from the photometric likelihood, functioning purely as a non-differentiable attention mask for the spatial-temporal constraints.
@@
Total Loss. The final objective function combines the uncertainty-weighted likelihood (data fidelity), the confidence-aware priors (regularization), and auxiliary perceptual constraints:
@@
where ensures perceptual quality using standard metrics (Wang et al., 2004; Zhang et al., 2018). Additionally, imposes a global silhouette constraint, aligning the accumulated opacity with the projected SMPL mask to guide coarse geometry. balances the contribution of each term.
@@
| B | + Uncert. Modeling | 24.15 | 0.9575 | 35.20 |
| C | + Spa. Con. () | 24.48 | 0.9592 | 32.50 |
| D | + Temp. Con. () | 24.62 | 0.9606 | 31.72 |
5. Experiments
5.1. Experimental Setup
Datasets. We evaluate U-4DGS on two standard benchmarks representing both synthetic and real-world occlusion scenarios. First, we utilize the ZJU-MoCap dataset (Peng et al., 2021b), which contains multi-view sequences of 6 subjects. Following the protocol of OccNeRF (Xiang et al., 2023), we simulate occlusions by masking the central 50% of the human region in the first 80% of the frames. We adopt a monocular setting for training, using only Camera 1 with 100 frames subsampled at an interval of 5, while the remaining 22 cameras are reserved for novel view evaluation. Second, to assess robustness in natural environments, we employ the OcMotion dataset (Huang et al., 2022), which features varying degrees of real-world human-object interactions. Consistent with prior works (Xiang et al., 2025; Sun et al., 2024), we select 6 representative sequences and train on sparse sub-sequences of 50 frames to challenge the model’s capability under limited observations.
Baselines. We benchmark our method against comprehensive state-of-the-art approaches, categorized into two distinct groups: (1) Standard Human Rendering. To demonstrate the impact of occlusion on conventional pipelines, we evaluate HumanNeRF (Weng et al., 2022), GaussianAvatar (Hu et al., 2024a), and GauHuman (Hu et al., 2024b). These methods are trained directly on the occluded sequences without specific handling mechanisms. (2) Occlusion-Aware Approaches. We compare against representative methods covering three mainstream technical paradigms: a) Scene Decoupling: OccNeRF (Xiang et al., 2023) and Wild2Avatar (Xiang et al., 2025), which utilize implicit fields or layer-wise parameterization to separate the human subject from environmental obstacles. b) Geometric Heuristics: OccGaussian (Ye et al., 2025) and SymGaussian (Jiang et al., 2025), which rely on rigid priors such as feature aggregation or left-right symmetry to infer missing geometry. c) Generative Priors: OccFusion (Sun et al., 2024) and GTU (Lee et al., 2024), which leverage 2D diffusion models to hallucinate the unobserved regions. For a fair comparison, all methods are evaluated using identical segmentation masks and SMPL pose priors.
Implementation Details: Our framework is implemented in PyTorch (Paszke et al., 2019) and trained on a single NVIDIA A100 PCIe GPU. The Probabilistic Deformation Network is instantiated as an 8-layer MLP (hidden dimension 256) with a skip connection at the 4-th layer. To capture high-frequency spatial-temporal dynamics, we apply sinusoidal positional encoding to input coordinates and time with frequency bands and , respectively.
We optimize the model for 10k iterations using the Adam optimizer (Kingma and Ba, 2015). Learning rates for canonical Gaussian attributes follow the schedule of GauHuman (Hu et al., 2024b), while the deformation network utilizes a rate of . Crucially, to preclude trivial solutions (where the model simply predicts infinite uncertainty to minimize the NLL), we adopt a progressive training strategy. For the initial 2k iterations, we deactivate the uncertainty branch and supervise geometry solely via standard loss. Subsequently, we activate the uncertainty head and switch to the MAP objective (Eq. 7) to jointly optimize for reconstruction fidelity and reliability. The regularization weights are set as: , , and .
5.2. Comparisons with State-of-the-Arts
Quantitative Results. Tab. 4.2 reports the performance on ZJU-MoCap and OcMotion benchmarks. On the synthetic ZJU-MoCap dataset, U-4DGS sets a new SOTA, achieving a PSNR of 24.62 dB. Notably, it surpasses the leading generative baseline, OccFusion (23.96 dB), validating that our uncertainty-guided aggregation yields superior fidelity without relying on potentially unstable diffusion priors. On the real-world OcMotion dataset, U-4DGS demonstrates exceptional robustness. Since the ground-truth of human appearance for occluded regions in real-world scenarios is unavailable, we calculate the metrics on visible pixels only. U-4DGS outperforms the runner-up by a substantial margin of 1.83 dB. Unlike standard rendering or scene decoupling methods that degrade under occlusions, U-4DGS effectively leverages Confidence-Aware Regularizations to shield the canonical geometry, proving that modeling heteroscedastic noise is decisive for handling unconstrained interactions.
Qualitative Results. Fig. 3 visualizes the rendering quality. Standard methods like GauHuman (GH) (Hu et al., 2024b) lack occlusion awareness, erroneously fusing environmental obstacles (e.g., bars or boxes) directly onto the human body. While generative approaches like OccFusion (OF) (Sun et al., 2024) can remove occluders, they suffer from inherent stochasticity, often yielding over-smoothed textures or hallucinations inconsistent with the subject’s identity. In contrast, U-4DGS achieves superior fidelity. By functioning as an adaptive gradient modulator, our uncertainty mechanism effectively attenuates gradients from artifacts. By strictly aggregating valid temporal information, our method faithfully restores clean, sharp, and physically consistent avatars, even under severe occlusions.
5.3. Ablation and Analysis
Ablation Studies. We validate the efficacy of each component by incrementally integrating it into the baseline to instantiate our complete MAP framework. Quantitative metrics on ZJU-MoCap are reported in Tab. 2, and visual comparisons are presented in Fig. 4. (1) Exp. A (Baseline): We establish a deterministic baseline OccGauHuman, structurally inheriting the pipeline from GauHuman (Hu et al., 2024b). It employs a standard deformation field optimized via loss, effectively assuming homoscedastic observation noise. As shown in Fig. 4 (Exp. A), limited by this rigid deterministic fitting, the model fails to distinguish between the human body and environmental obstacles. Consequently, it indiscriminately minimizes geometric error, leading to severe artifact adhesion at occlusion boundaries. (2) Exp. B (+ Uncertainty Modeling): Replacing the deterministic field with our Probabilistic Deformation Network and enabling Joint Rasterization yields a substantial quality leap. This step introduces the Likelihood term of our MAP objective. By functioning as an adaptive gradient modulator, the learned uncertainty effectively attenuates gradients from unreliable observations, successfully removing occlusion artifacts. However, without regularization priors, the geometry in regions lacking reliable visual cues remains noisy and ill-defined (see Exp. B), as the probabilistic loss solely down-weights error without explicitly guiding completion. (3) Exp. C (+ ): The introduction of Confidence-Guided Spatial Consistency imposes the first Prior term. By leveraging uncertainty to selectively propagate spatial validity from confident neighbors, this constraint effectively regularizes the noisy geometry, restoring a coherent and complete body shape. (4) Exp. D (+ ): Finally, incorporating Uncertainty-Weighted Temporal Inertia completes the U-4DGS framework. This temporal prior ensures trajectory smoothness and further refines surface details, achieving the most faithful reconstruction with high physical plausibility.
@@ 5.3. Ablation and Analysis
Visualization of Uncertainty Maps. To validate the efficacy of our heteroscedastic noise modeling, we visualize the rendered uncertainty map in Fig. 5. In our MAP framework, denotes the estimated scale of observation noise and is tasked with autonomously distinguishing valid signals from occlusion artifacts. As evident in Fig. 5, the network assigns low uncertainty (depicted in dark purple) to visible body parts, allowing the photometric supervision to drive the optimization in these trusted regions. Crucially, for anatomical regions that are occluded in the current frame but belong to the canonical topology, the network autonomously predicts significantly high uncertainty (depicted in bright yellow). This visualization empirically confirms that functions as an adaptive gradient modulator. By automatically increasing the denominator in the NLL objective (Eq. 7), it effectively attenuates gradients in these regions lacking reliable visual cues, preventing the canonical geometry from adhering to environmental occlusions.
[…]
| OccNeRF | 20 h | 0.20 |
| OccGaussian | 6 min | 169 |
| OccFusion | 10 min | 151 |
| U-4DGS (Ours) | 7.5 min | 155 |
Efficiency Analysis. We present a quantitative comparison of training time and rendering speed (FPS) in Table 3. Implicit methods such as OccNeRF are severely bottlenecked by volumetric ray-marching, resulting in prohibitive training times and non-real-time inference. Generative prior-based approaches like OccFusion achieve real-time rendering but require relatively longer training (10 minutes) due to the integration of diffusion models. In contrast, U-4DGS preserves the efficiency inherent to 3D Gaussian Splatting, achieving an interactive rendering speed of 155 FPS. Furthermore, the introduction of the Probabilistic Deformation Network and the joint rasterization pipeline introduces minimal computational overhead compared to heuristic-based methods (e.g., OccGaussian), requiring only 7.5 minutes for training.
6. Conclusion
@@
In this paper, we present U-4DGS, a robust framework that reformulates monocular human rendering as an MAP estimation problem to handle severe environmental occlusion. Departing from paradigms that rely on stochastic hallucinations or rigid priors, we explicitly model the heteroscedastic nature of observation noise. By integrating a Probabilistic Deformation Network with a Joint Rasterization pipeline, our method establishes an adaptive gradient modulator, enabling the optimization to autonomously distinguish between valid geometric cues and occlusion artifacts. Furthermore, we leverage the learned uncertainty to drive Confidence-Aware Regularizations, preventing geometric drift and ensuring structural integrity in regions lacking reliable visual cues. Extensive evaluations on ZJU-MoCap and OcMotion demonstrate that U-4DGS significantly outperforms state-of-the-art approaches, achieving superior robustness and photorealism.